### v1

@@ I Introduction
Up to now, there has been only two neutron detectors: LAND bib:Pawlowski2023 and NeuLAND bib:Boretzky2021 , both developed and constructed at GSI, which are able to measure neutrons in heavy ion collisions at beam energies below 1A GeV. The LAND consists of 10 mutually perpendicular layers with 20 individual long scintillation detectors in each layer. Each individual detector consists of 10 alternating layers of plastic scintillation and iron plates with a thickness of 0.5 cm and has a size of $10\times 10\times 200$ cm3, the light from two ends of which is read by two photomultiplier tubes. The full geometric size of the LAND detector is $200\times 200\times 100$ cm3. A typical resolution of time measurements with the LAND is reported to be 250 ps.
A new large area neutron detector NeuLAND bib:Boretzky2021 has been designed and partially constructed to study the properties of nuclear matter produced in collisions of radioactive nuclei, within the R3B project at the FAIR accelerator facility in Darmstadt, Germany. In contrast to LAND, NeuLAND consists only of active layers of plastic scintillation detectors with a size of $250\times 5\times 5$ cm3 each. The light from the opposite ends of the detectors is read by two photomultiplier tubes. NeuLAND is composed of 60 active layers with 50 scintillation detectors in each layer and has a total size of $250\times 250\times 300$ cm3. The time resolution of NeuLAND is estimated to be 150 ps, the spatial resolution of the first interaction point of the neutron is about 1.5 cm, and the efficiency of detecting each neutron in an event is better than 95% for kinetic energies of 400 – 1000 MeV.
Taking into account the significantly higher energy (up to 4 GeV) of the neutrons produced in nucleus-nucleus collisions in the BM@N experiment and the intense development of hadron showers in the detector volume at these energies, it is impossible to use large-size neutron detectors like LAND and NeuLAND. Therefore, a new concept of Highly Granular Neutron Detector (HGND) has been proposed. Instead of the long scintillator plates used in the LAND and NeuLAND, it is proposed to use the highly granular transverse structure of the active elements of the detector to identify neutrons and measure their energy with the time-of-flight method with good resolution. At the same time, the multilayer longitudinal sampling makes it possible to achieve a high efficiency of neutron detection bib:Guber2024 . It was proposed to build the active layers of the HGND as ($11\times 1$ 11) arrays of small scintillator cells with individual light readout by silicon photomultipliers (SiPM). Each cell is expected to provide a time resolution of $\sim$150 ps. Copper absorbers placed between the active layers will provide a rather high neutron interaction probability of the HGND.
@@ II.1 Concept of the HGND
The original concept of the full-scale HGND consisting of 16 alternating active layers of 121 plastic scintillation detectors grouped in the $11\times 1$ 11 array, with copper absorber plates in between, has been presented in detail in bib:Guber2024 . The absorber plates (44 $\times$ 44 $\times$ 3 cm3) and the scintillator layers (44 $\times$ 44 $\times$ 2.5 cm3) will be mounted on a common support frame. The total length of the HGND is about 1 m and corresponds to $\sim$3 nuclear interaction lengths $\lambda_{\rm int}$. The first scintillation layer of the HGND will be used as a VETO detector for charged particles. The scintillator layers will be assembled from 4 $\times$ 4 $\times$ 2.5 cm3 plastic scintillators (cells) based on polystyrene with additions of 1.5% paraterphenyl and 0.01% POPOP produced at JINR. The light read-out from the center of entrance surface of each cell is conducted with EQR15 11-6060D-S silicon photomultipliers bib:Eqr15_2025 placed on the printed circuit board (PCB) that hosts preamplifier and Low-Voltage Differential Signal (LVDS) comparators, generating Time over Threshold (ToT) bib:KARPUSHKIN2024169739 signals for the read-out schematics. More technical details of the HGND design as well as results of the simulations can be found in bib:Guber2024 .
@@ II.2 Design of the HGND prototype
Each of the scintillator layers consists of a 3 $\times$ 3 matrix of individual scintillators with a transverse area of 4 $\times$ 4 cm2 and a thickness of 2.5 cm. The light from each of the nine scintillators is read out by an individual MPPC Hamamatsu S13360-6050PE bib:Hamamatsu2025mppc with a sensitive area of 6 $\times$ 6 mm2 mounted on a four-layer PCB with readout electronics, necessary connectors and a sensor for temperature correction. Each of the nine readout channels consists of a photodetector, an input amplifier, a buffer amplifier with 50 $\Omega$ line capability. Every PCB has voltage regulators for the positive and negative supply for the amplifiers. This board is attached directly to the scintillators in the layer. The absorber, scintillators and the PCB with photodetectors and electronics are packaged in a light-protective module made by 3D printing.
@@ III.2 Reconstruction of neutron kinetic energy
The measured probability distribution to trigger a given scintillator layer by particles with the highest velocity in the event is presented in Fig. 5 as a function of the layer number. As can be seen, neutrons emitted both in hadronic collisions and EMD interact mostly after the 7th layer. Photons which hit the HGND prototype were rejected by imposing the condition that if the first layer after the VETO layer is triggered, the event is discarded as associated with a photon. Due to such a selection, about 20% of the events of hadronic interactions as well as of EMD were rejected. The systematic uncertainty of the photon rejection procedure contributes to the total systematic uncertainty of these measurements.
@@ IV Comparison with Monte Carlo simulations
The selection procedure applied to the detected events and described in Sec. III.1 was used also in modeling. The efficiency calculated with Geant4 bib:Geant4 for the detection of a single neutron of a given energy in the HGND prototype is shown in Fig. 7. As can be seen, due to more developed hadronic showers induced by more energetic neutrons, the detection efficiency is higher at higher energies. A polynomial fit to the modeled efficiency, also shown in Fig. 7, was used to correct the number of reconstructed neutrons depending on their energy.
@@
The detection of neutrons by the HGND prototype in its forward ($0.7^{\circ}$) position was modeled with the BmnRoot software taking into account the detailed geometry of the BM@N experiment. In these simulations, the time resolution of the scintillation cells was set to 270 ps. We considered hadronic collisions only with the centrality of 0-60% as selected in the experiment by the CCT trigger, and applied the corresponding centrality selection to simulated events. In particular, events with an impact parameter of less than 9.54 fm ($<$60%) were modeled.
The correlations between the emission angle $\Theta$ and neutron kinetic energy $T_{n}$ obtained in simulations with the DCM-QGSM-SMM and RELDIS models are presented in Fig. 8. As seen, all neutrons from EMD events are characterized by a very narrow angular distribution, and the average neutron energy is equal to the beam energy of 3.8 GeV. Basically, the energy distribution of EMD neutrons is also very narrow: the calculated energies of all EMD neutrons are between 3.4 and 4.2 GeV, within ±$\pm 10$% of the average bib:Pshenichnov2024 .
Figure 8: Correlations between the emission angle $\Theta$ and kinetic energy $T_{n}$ of neutrons generated in 124Xe–CsI collisions. Results of the DCM-QGSM-SMM model (0-60% centrality, left) and RELDIS model (right) are shown.
@@
DCM-QGSM-SMM 3.37 $\pm$ 0.02 39.41 $\pm$ 0.20 1.33 $\pm$ 0.01
RELDIS 34.21 $\pm$ 0.25 61.49 $\pm$ 0.45 20.38 $\pm$ 0.15
While the HGND prototype is $\sim 50$% more efficient in detecting more forward-focused EMD neutrons which hit its forward surface with respect to neutrons from hadronic fragmentation, there is a drastic difference in the geometrical acceptance of neutrons from these two processes (3.37% vs 34.21%). Taking into account the resulting $acc\times\varepsilon$, one can conclude that the HGND prototype is much more efficient in detecting neutrons from EMD in comparison to the detection of forward neutrons from hadronic collisions.
One can note that the detection efficiency (61.5%) calculated for EMD neutrons with $\langle T_{n}\rangle=3.8$ GeV is slightly lower than the efficiency (66%) calculated for monoenergetic neutrons with $T_{n}=3.8$ GeV, see Fig. 7. The difference in calculated efficiency between EMD and hadronic interactions is due to the $\sim 1.5$ times different average multiplicity of neutrons hitting the detector. Since it is impossible to reconstruct more than one neutron in an event with the current detector configuration, the probability of detecting all neutrons in the event is essentially reduced.

### v2

@@ I Introduction
Up to now, only two neutron detectors have been developed: LAND bib:Pawlowski2023 and NeuLAND bib:Boretzky2021 . Both were developed and constructed at GSI in Darmstadt, Germany to measure neutrons from heavy-ion collisions at beam energies below 1A GeV. The LAND consists of 10 mutually perpendicular layers with 20 individual long scintillation detectors in each layer. Each individual detector consists of 10 alternating layers of plastic scintillation and iron plates with a thickness of 0.5 cm and has a size of $10\times 10\times 200$ cm3, the light from two ends of which is read by two photomultiplier tubes. The full geometric size of the LAND detector is $200\times 200\times 100$ cm3. A typical resolution of time measurements with the LAND is reported to be 250 ps.
A new large area neutron detector NeuLAND bib:Boretzky2021 has been designed and partially constructed to study the properties of nuclear matter produced in collisions of radioactive nuclei, within the R3B project at the FAIR accelerator facility at GSI. In contrast to LAND, NeuLAND consists only of active layers of plastic scintillation detectors with a size of $250\times 5\times 5$ cm3 each. The light from the opposite ends of the detectors is read by two photomultiplier tubes. NeuLAND is composed of 60 active layers with 50 scintillation detectors in each layer and has a total size of $250\times 250\times 300$ cm3. The time resolution of NeuLAND is estimated to be 150 ps, the spatial resolution of the first interaction point of the neutron is about 1.5 cm, and the efficiency of detecting each neutron in an event is better than 95% for kinetic energies of 400 – 1000 MeV.
Taking into account the significantly higher energy (up to 4 GeV) of the neutrons produced in nucleus-nucleus collisions in the BM@N experiment and the intense development of hadron showers in the detector volume at these energies, it is impossible to use large-size neutron detectors like LAND and NeuLAND. Therefore, a new concept of Highly Granular Neutron Detector (HGND) has been proposed. Instead of the long scintillator plates used in the LAND and NeuLAND, it is proposed to use the highly granular transverse structure of the active elements of the detector to identify neutrons and measure their energy with the time-of-flight method with good resolution. At the same time, the multilayer longitudinal sampling makes it possible to achieve a high efficiency of neutron detection bib:Guber2024 . It was proposed to build the active layers of the HGND as ($11\times 1$ 11) arrays of small scintillator cells with individual light readout by silicon photomultipliers (SiPM). Each cell is expected to provide a time resolution of $\sim$150 ps. Copper absorbers placed between the active layers will provide a rather high probability of neutron interaction with the HGND.
@@ II.1 Concept of the HGND
The original concept of the full-scale HGND consisting of 16 alternating active layers of 121 plastic scintillation detectors grouped in the $11\times 1$ 11 array, with copper absorber plates in between, has been presented in detail in bib:Guber2024 . The absorber plates (44 $\times$ 44 $\times$ 3 cm3) and the scintillator layers (44 $\times$ 44 $\times$ 2.5 cm3) will be mounted on a common support frame. The total length of the HGND is about 1 m and corresponds to $\sim$3 nuclear interaction lengths $\lambda_{\rm int}$. The first scintillation layer of the HGND will be used as a VETO detector for charged particles. The scintillator layers will be assembled from 4 $\times$ 4 $\times$ 2.5 cm3 plastic scintillators (cells) based on polystyrene with additions of 1.5% paraterphenyl and 0.01% POPOP produced at JINR. The light read-out from the center of entrance surface of each cell is conducted with EQR15 11-6060D-S silicon photomultipliers bib:Eqr15_2025 ; bib:Guber2024b placed on the printed circuit board (PCB) that hosts preamplifier and Low-Voltage Differential Signal (LVDS) comparators, generating Time over Threshold (ToT) bib:KARPUSHKIN2024169739 signals for the read-out schematics. The methods to identify neutrons in high multiplicity neutron events in the presence of the background from charged particles and photons are currently under development. More technical details of the HGND design as well as results of the simulations can be found in bib:Guber2024 .
@@ II.2 Design of the HGND prototype
Each of the scintillator layers consists of a 3 $\times$ 3 matrix of individual scintillators with a transverse area of 4 $\times$ 4 cm2 and a thickness of 2.5 cm. Light from each of the nine scintillators is read out by an individual SiPM (Hamamatsu S13360-6050PE) bib:Hamamatsu2025mppc with a sensitive area of 6 $\times$ 6 mm2, mounted on a four-layer PCB with readout electronics, necessary connectors and a sensor for temperature correction. These SiPMs were used in the HGND prototype due to the limited availability of EQR15-11-6060D-S units at the time of prototype construction. However, the full-scale HGND will adopt EQR15-11-6060D-S SiPMs due to their lower cost and superior timing resolution, as demonstrated in electron-beam tests bib:Guber2024b . Each of the nine readout channels consists of a photodetector, an input amplifier, a buffer amplifier with 50 $\Omega$ line capability. Every PCB has voltage regulators for the positive and negative supply for the amplifiers. This board is attached directly to the scintillators in the layer. The absorber, scintillators and the PCB with photodetectors and electronics are packaged in a light-protective module made by 3D printing.
@@ III.2 Reconstruction of neutron kinetic energy
The measured probability distribution to trigger a given scintillator layer by particles with the highest velocity in the event is presented in Fig. 5 as a function of the layer number. As can be seen, neutrons emitted both in hadronic collisions and EMD interact mostly after the 7th layer. Photons which hit the HGND prototype were rejected by imposing the condition that if the first layer after the VETO layer is triggered, the event is discarded as associated with a photon. Due to such a selection, about 20% of the events of hadronic interactions as well as of EMD were rejected. The uncertainty of the photon rejection procedure is expected to be one of the main sources of systematic uncertainties in measuring neutron yields. It was estimated from the change in neutron yield obtained without photon rejection with respect to the standard reconstruction. Other contributions to the systematic uncertainty of the measurements of neutron yield and neutron flow in the BM@N experiment were not detailed in this work. This will be a subject of future studies.
@@ IV Comparison with Monte Carlo simulations
The selection procedure applied to the detected events and described in Sec. III.1 was used also in modeling. The efficiency calculated with Geant4 bib:Geant4 for the detection of a single neutron of a given energy in the HGND prototype is shown in Fig. 7. In this modeling neutrons were transported directly to the front surface of the detector placed in vacuum and were not accompanied by any other particles. As can be seen, due to more developed hadronic showers induced by more energetic neutrons, the detection efficiency is higher at higher energies. A polynomial fit to the modeled efficiency, also shown in Fig. 7, was used to correct the number of reconstructed neutrons depending on their energy.
@@
The detection of neutrons by the HGND prototype in its forward ($0.7^{\circ}$) position was modeled with the BmnRoot software taking into account the detailed geometry of the BM@N experiment. In these simulations, the time resolution of the scintillation cells was set to 270 ps. We considered hadronic collisions only with the centrality of 0-60% as selected in the experiment by the CCT trigger, and applied the corresponding centrality selection to simulated events. In particular, events with an impact parameter of less than 9.54 fm were selected from event files generated with the DCM-QGSM-SMM model to represent 0–60% centrality. As demonstrated in Ref. bib:Segal2023 , the results of centrality estimators based on the multiplicity of charged particles and on signals from forward calorimeters (which are impacted by forward neutrons, among other particles and nuclear fragments) are consistent within their uncertainties. In particular, all methods associate the interval $[0,b_{max}]$, where $b_{max}$ $\sim$ 9$\pm$1 fm, with the centrality range of 0–60%. Therefore, the result of Ref. bib:Segal2023 obtained for 4A GeV Xe+CsI collisions justifies the selection of b $<$ 9.54 fm in the DCM-QGSM-SMM model.
The correlations between the emission angle $\Theta$ and neutron kinetic energy $T_{\rm n}$ in the laboratory system obtained in simulations with the DCM-QGSM-SMM and RELDIS models are presented in Fig. 8. The spread of kinetic energy $T_{\rm n}$ is essential in hadronic fragmentation events. For example, few percent of forward neutrons from hadronic fragmentation of 124Xe have kinetic energy up to 5.5 GeV, i.e., much higher than the nominal beam kinetic energy of $T_{\rm lab}=3.8$ GeV/nucleon. In a seminal paper on nuclear fragmentation by A.S. Goldhaber bib:GOLDHABER1974306 , the data on fragment momentum distributions resulting from peripheral nucleus-nucleus collisions were parameterized as $\exp(-p^{2}/2\sigma^{2})$ with a characteristic width of the distribution of $\sigma$ = 90–100 MeV/c estimated for fragmentation nucleons. A theoretical estimation $\sigma^{2}=\langle p^{2}\rangle/3=p^{2}_{F}/5$ has been obtained in Ref. bib:GOLDHABER1974306 by taking randomly an individual nucleon released in fragmentation and considering its intranuclear Fermi motion. With a typical value of the Fermi momentum of $p_{F}$ = 230 MeV/c the value of $\sigma\sim$ 100 MeV/c is confirmed, see also Ref. bib:PhysRevC.65.051602 which reports $\sigma\sim$ 112–116 MeV/c. It can be expected that following ”the $3\sigma$ rule” almost all spectator neutrons have $p<300$ MeV/c ($T<46.8$ MeV) in the center of mass system of the initial 124Xe nucleus. In the extreme case of a neutron emitted in the beam direction with $p<~{}300$ MeV/c in the rest frame of a beam nucleus, the boost with a Lorentz factor of $\gamma$ = 5.0469 provides $T_{\rm n}=5.52$ GeV in the laboratory system, in a good agreement with the DCM-QGSM-SMM modeling.
Figure 8: Correlations between the emission angle $\Theta$ and kinetic energy $T_{\rm n}$ of neutrons generated in 124Xe–CsI collisions. Results of the DCM-QGSM-SMM model (0-60% centrality, left) and RELDIS model (right) are shown.
As seen in Fig. 8, all neutrons in simulated EMD events are characterized by a very narrow angular distribution, and the average neutron energy is equal to the beam energy of 3.8 GeV. Basically, the energy distribution of EMD neutrons is also very narrow: the calculated energies of all EMD neutrons are between 3.4 and 4.2 GeV, within ±$\pm 10$% of the average bib:Pshenichnov2024 . Since the neutrons emitted by electromagnetically excited 124Xe nuclei mostly result from decays of the low-lying Giant Dipole Resonance in 124Xe, their kinetic energies in the rest frame of a beam nucleus are below few MeV.
@@
DCM-QGSM-SMM 3.37 $\pm$ 0.02 39.69 $\pm$ 0.20 1.34 $\pm$ 0.01
RELDIS 36.77 $\pm$ 0.27 55.77 $\pm$ 0.41 20.51 $\pm$ 0.15
While the HGND prototype is $\sim 50$% more efficient in detecting more forward-focused EMD neutrons which hit its forward surface with respect to neutrons from hadronic fragmentation, there is a drastic difference in the geometrical acceptance of neutrons from these two processes (3.37% vs 36.77%). Taking into account the resulting $acc\times\varepsilon$, one can conclude that the HGND prototype is much more efficient in detecting neutrons from EMD in comparison to the detection of forward neutrons from hadronic collisions.
It should be noted that the detection efficiency (55.8%) calculated for EMD neutrons with $\langle T_{\rm n}\rangle=3.8$ GeV using the realistic geometry of the BM@N setup is lower than the efficiency (66%) calculated for monoenergetic neutrons with $T_{\rm n}=3.8$ GeV shot directly from a particle source to the front surface of the HGND prototype, see Fig. 7. On average, approximately 1.02 neutrons hit the front detector surface per EMD event. Therefore, in some 2% of EMD events, a pair of neutrons hits the detector surface. Typically, the second neutron is not detected in these events due to the small transverse size of the HGND prototype and the employed reconstruction algorithm. In addition, in the realistic geometry, some neutrons are accompanied by charged particles or photons resulting from the secondary fragmentation of 124Xe in the structural components of the BM@N setup or in the air. As described in Sec. III.2, these neutrons are also rejected due to the photon rejection procedure. Taken together, these factors explain the aforementioned difference (55.8% vs 66%).
The difference in efficiency values calculated for EMD and hadronic interactions is due to the $\sim 1.5$ times different average multiplicity of neutrons hitting the detector. Since it is impossible to reconstruct more than one neutron in an event with the current detector configuration, the probability of detecting all neutrons in the event is essentially reduced.
@@
Figure 11: Reconstructed kinetic energy spectra of neutrons from hadronic fragmentation (left) and EMD (right) of 3.8A GeV 124Xe on CsI target obtained in simulations with three different settings for time resolution: 0, 200 and 270 ps.
@@
While the extension of neutron kinetic energy spectra up to 5.5 GeV in hadronic events is the effect of the relativistic kinematics, a further extension of the reconstructed energy spectra up to 10 GeV is caused by a limited resolution of the time-of-flight measurements with the HGND prototype. The neutron kinetic energy spectra reconstructed by the time-of-flight method using the fastest trigger of the scintillation cell are shown in Figs. 6 and 10, and they were obtained assuming the time resolution of 270 ps. In order to demonstrate the sensitivity of reconstruction results to the choice of time resolution, the reconstructed spectra were obtained also with the time resolution parameters of 0 and 200 ps in addition to the nominal 270 ps and they are compared in Fig. 11. In this figure the impact of the time resolution on the shape of the reconstructed spectra is clearly visible. In the ideal case of 0 ps resolution the initial spectra from event generators were reproduced quite well, but the artifact of the extension of the reconstructed spectra up to 10 GeV is also seen. One can note that in the present work the extreme case of detecting forward most energetic neutrons has been considered and used for detector calibration. In future studies of neutron flow in the mid-rapidity region the full-scale HGND will be used to detect neutrons with $T_{\rm n}<3$ GeV. As seen in Fig. 11, the impact of the time resolution on the reconstructed spectra is essentially alleviated at $T_{\rm n}<3$ GeV.